A poll was taken of 500 College graduates. The college graduates sampled were asked whether Americans feel satisfied with their lives; 350 responded “very satisfied” Use this date to obtain a 97% confidence interval for the corresponding population proporation, p.
Answer by ewatrrr(10682)

 
Hi
A poll was taken of 500 College graduates.
The college graduates sampled were asked whether Americans feel satisfied with their lives;
350 responded “very satisfied” Use this date to obtain a 97% confidence
interval for the corresponding population proporation, p.
p-hat = 350/500 = .7
97% confidence then z = 2.1701
ME = 2.1701*sqrt[0.7*0.3/500] = 0.04447
97%CI: 0.7-0.04447 < p < 0.7+0.04447
97%CI: 0.6555< p < 0.7448
Summary of other critical regions
a a/2 crtical regions
90% 0.1 5% z <-1.645 z >+1.645
95% 0.05 2.50% z <-1.96 z >+1.96
98% 0.02 1% z <-2.33 z >+2.33
99% 0.01 0.50% z<-2.575 z >+2.575